William's uncle is three times as old as William. In 12 years, his uncle will be twice h
is age. How old is William now.

Answer by stanbon(75887)

William's uncle is three times as old as William. In 12 years, his uncle will be twice his age. How old is William now.
-----------
Equation:
u = 3w
u+12 = 2(w+12)
-----
Substitute for "u" and solve for "w":
3w+12 = 2w+24
w = 12 yrs. (William's age now)
====================================
cheers,
Stan H.